Seismic activity has increased since yesterday. Three (3) earthquakes were located. Two events occurred under the southern end of St George's Hill at depths of 4-6km. One event was also located off the west coast of Montserrat. The number of small, long period events increased this afternoon. These events were slightly larger and more frequent than has been observed over the past few days. This suggests that slow dome growth is still occurring. Larger long period events have also continued throughout the period at a similar rate to that observed yesterday. Broadband tremor has been less dominant today.
General maintenance of the seismometer network has been carried out today.
Persistent rain and low cloud prevented any EDM measurements today. The GPS sites at St George's Hill, Radio Antilles and Harris' Lookout were occupied today, but the data are still being processed. Data collected over the past two days has been processed. No changes were observed from the previous occupations of the sites.
The poor weather conditions prevented any visual observation of the volcano.